OpenStudy (anonymous):

In EFG, what is the approximate length of if the measure of angle G is 50° and EF is 12 cm?

OpenStudy (anonymous):

EF being the opposite side of angle G

OpenStudy (anonymous):

of what?? is this all the given?

OpenStudy (anonymous):

Yes an EFG triangle is given with F being a right angle, G being 50 degrees, and line EG being adjacent to G. EF is opposite.

OpenStudy (anonymous):

right triangle so apply trigonometry.sin(50)=12/EG..now use your calculator

OpenStudy (anonymous):

The only problem is that my calculator isn't capable of this type of math. And they also didn't really show me how to do this on a calculator.

OpenStudy (anonymous):

windows has a calculator with trig. fcts...

OpenStudy (anonymous):

Okay, how would I enter this into the calculator?

OpenStudy (anonymous):

sin(50)=12/EG=>EG=12/sin(50) so 12/50 sin

OpenStudy (anonymous):

I typed all that in and received no answer. Do you mind telling me the answer? This is the only question I am having trouble with

OpenStudy (anonymous):

??